# Counterfactual Effect Generalization: A Combinatorial Definition

### v4

@@ 4.1 Randomization and Causal effects
We will use the notation often employed in Structurual Causal Models (SCM) to relate the previous concepts to current research. In SCM models, variables are typically specified by an unknown function of a set of endogeneous variables and an exogeneous error. We have, similarly, specified DGPs $Z=\{X,U\}$ with a collection of exogenous factors $U$, in addition to the observed $X$. Factors in $U$ are unobserved and not affected by $X$, but affect $X$. More specifically, a SCM specifies a collection of functions $f_{a}$ indexed by endogenous variables, $f=(f_{a}:a,b,c,...,[m])$. The SCM also calls for the specification of the relation between random variables $U$ and $X$. This is done through a set of functions $f=(f_{W},f_{a},f_{y})$, where $W$ are the other (non-treated) in-sample factors, $W=X-\{a\}$, and $y$ is the outcome of interest. The values of $W$, $A$, and $Y$ are then deterministically assigned by $U$, according to  [van der Laan and Rose, 2011]
@@
Randomized controlled trials (RCTs) are one type of controlled experiment where subjects are randomized to receive a specific treatment. For example, if each subject was assigned to one of two levels of treatment based on the flip of a coin. Then the differences between the two groups would be solely due to treatment, as all other factors would be balanced (free os selection bias), up to random error. In an RCT the treatment assignment process, $P(A=1|W)$, is known (e.g., constant and $0.5$ as in a coin flip). This is very different from non-experimental data where treatment, or treatment exposure, is not assigned. In SCM notation, a RCT can be, nonetheless, written as in Eq.(15-16,18),
@@
The only difference of Eq.(14) from Eq.(15-16,18) is the structural equation for $f_{a}$. The baseline variables $W$ play no role in the generation of $A$ in Eq.(16). These alternative assumptions can be briefly illustrated with causal graphs, Fig.2(c). Since $a$ is randomized, this implies that $U_{a}$ is independent of $U_{y}$ and $U_{W}$, and, in this case, no arrows connect $U_{a}$ to $U_{y}$ and $U_{a}$ to $U_{w}$.
4.2 Effect Background Randomization
Let $p=\{p_{1},p_{2},...,p_{m}\}$ be a sample gap background assignment. Consider a set of subsequent unobserved $U=Z\backslash X$ backgrounds $\Pi=\{\pi_{0},\pi_{1},\pi_{2},...\}$. Generalizability requires effects to be valid across any $\Pi$, irrespective of $p$. Eq.(8) stipulated two gap distributions that could allow us to estimate effects despite unobserved confounding. Imagine we keep $p$ constant from a time $t=0$ to $\delta t$ while allowing $\Pi_{n}(Z/X)$ to change. Can we say anything about whether the backgrounds $\pi_{0}$ and $\pi_{\delta t}$ are random (i.e., are as separated as in a uniform distribution), $\pi_{0}\perp\pi_{t}$? Fix $p$ and two backgrounds $\pi_{0},\pi_{1}\in\Pi_{n}(Z/X)$ separated by only one inversion. The likelihood of reaching $\pi_{1}$ from $\pi_{0}$ after one time step is high in this case. With increasing inversions in $\pi_{0}$, Eq.(11), there should be some period $\delta t$ where that likelihood becomes comparable to that of a uniform distribution - where $\binom{m}{2}$ inversions are generated at each time. Eq.(11) formulates that relationship. With an increasing $t$ (and $p$ with small $s$), this happens quickly, as its right-most term converges quickly to $1$. The choice of a background $\pi_{1}$ from $\pi_{0}$ is then as likely as any other background. Fig.2(b) shows this limit for the simulations in Sect.6 Experiments, which include correlated and non-stationary cases. The variation separation distance (y-axis) is a measure of randomness (i.e., formulated as a distance to an uniform distribution)  [Bayer and Diaconis, 1992, Diaconis, 1996]. The panel shows the distance between a fixed background $\pi_{0}$ and the subsequent in samples where $p$ is fixed and $s<0.25$ (these are exhaustively enumerated using the tests and procedures described next, distances where normalized to their per-factor range for display). The upper panel shows the same for other samples $s>0.75$. These combinatorial considerations suggest limits where some guarantees on effect background randomization can be made. Notice that $\binom{m}{2}$ in Eq.(11) is the number of effects observation in a single square. We will therefore use, more specifically, the number of inversions in squares to define testable conditions for background randomization in the next section.
[…]
The previous discussion indicates that, with random assignment of backgrounds to fixed treatments, as the sample size ($n$) approaches infinity, by the law of large numbers, the average characteristics of the treatment and control groups, $P(Z-\{a\}<a)$ and $P(Z-\{a\}>a$, converge to population averages. For a single factor $a$, we visualized these two groups as the upper and lower triangles of sample squares (i.e., in subsamples fulfilling the previous conditions). This can be written as
@@ 4.2 Effect Background Randomization
With convergence, the potential bias introduced by correlations in unobserved or uncontrolled factors ($U$) is mitigated. On average, any unobserved or uncontrolled factors that could potentially bias effect estimates will be equally distributed between the two effect observation orders, as required by Principle.1. In a square, this is true for all in-sample variables simultaneously ${\rm I\kern-1.99997ptE}\{\textrm{Var}[\,\Delta y(a)\,]\}\approx 0$. Notice that according to Eq.(19) this also implies that variables are expected to be confounded in equal amounts in this limit.
Proposition 3.
With random unobserved effect backgrounds and sufficiently large sample sizes $n$, subsample variables $a\in X$ converge asymptotically to true population treatment effects, despite $U{-}X$ correlations.
Availability of effect observations with these characteristics (and thus which strategy to use) is to be determined by statistical hypotheses tests. Hypotheses tests for these three conditions are similar but increasingly restrictive (i.e., requires larger samples).
4.3 Square Enumeration and Gap Inversion Tests
All three previous scenarios, and Principle.1, are related to permutation inversions. They were that $\pi_{0}$ changes, $\pi_{0}$ changes while $p$ stays constant for $\delta$ time, and $\pi_{0}$ changes while $p$ stays constant for $m\times\delta$ time. Inversions appear in common tests of i.i.d. sampling (’stationarity’) in ordinal data, such as in Kendall’s $\tau$ and Man-Whitney (when $X$ is binary) tests. Any pair of observations $(x_{i},y_{i})$ and $(x_{j},y_{j})$, where $i<j$, are said to be ’concordant’ if the sort order of $(x_{i},x_{j})$ and $(y_{i},y_{j})$ agrees: that is, if either both $x_{i}>x_{j}$ and $y_{i}>y_{j}$ holds or both $x_{i}<x_{j}$ and $y_{i}<y_{j}$; and be ’discordant’ otherwise. The number of discordant pairs between $x$ and $y$ is equal to the number of inversions that permutes the $y$-sequence into the same order of $x$. The statistic can also be seen as a generalised measure of a classifier’s separation power for more than two classes  [Hand and Till, 2001].
We formalize the previous conditions for non-experimental samples as any of their subsamples $C$, $C\subset X$, passing the following statistical hypothesis tests over inversions in sample ordinal and gap statistics,
[…]
Principle.1 is also related to inversions and says that confounders, causes and spurious factors have distinct effect variances under increasing inversions (e.g., the number of times confounder $b$ appears before its root cause). According to the principle, the observation of one inversion allows for the separation of a single cause-confounder pair (Appendix.A,B). We can generalize the previous conditions to non-experimental samples which are not sequentially sampled by enumerating squares for a specific sample unit $x_{0}$ using inter-unit combinatorial properties. In fully observed samples, $U=\emptyset$, square enumeration ensures the applicability of Principle.1. In samples with unobserved factors, $U\neq\emptyset$, enumeration with the additional conditions over gap inversions in Eq.(20) allows the same. Much like we can observe single factor differences in samples, and their effects, according to square enumeration, we can observe factor permutations and their effects, by combining single differences and their effects carefully. The panel in Fig.2(d) illustrates this process for two factors ($m=2$). It illustrates the differences (circles) and intersections (gray), from $x_{0}$, that characterize each cell in a square of size $2$. We say that by observing differences $a$, $b$ and $ab$ from $x_{0}$, we ’observed’ a permutation for these factors. That’s because these differences - $(x_{0}{-}x_{1})=\{a\}$, $(x_{0}{-}x_{2})=\{b\}$ and $(x_{0}{-}x_{3})=\{ab\}$ allow us to recover the effects of permuting factors $\{a,b\}$ (conditional on $x_{0}$), which are not directly observed in the sample. Notice that not all sample pairs with differences with $\{a\}$ and $\{ab\}$ would belong to the second square column, as there are also conditions on their intersection, $(x_{0}{-}x_{1})\cap(x_{0}{-}x_{3})=\{b\}$. It is the fulfillment of both combinatorial conditions that allow us to take them as (conditional) effect observations. The square full combinatorial structure generalizes this to $m>2$ and uses the $2^{m}$ $m$-way differences potentially available in samples. Further combinatorial, computational and statistical aspects of this enumeration process are described in Appendix.A,B,I.
Fig.1(a) summarizes this solution for effect estimation, where square enumeration ensures the inversions required for Principle.1 in in-sample variables, and tests over their their gaps, Eq.(20), in out-of-sample. For sets of units (and thus mutual effect observations) where these combinatorial and gap conditions hold simultaneously, variance in effects can become asymptotically indicative of the true status of factors as causes, Principle.1, Eq.(3) of their effects, and Eq.(4) of their effects generalizability. Variance becomes, in such conditions, a measurement of ’noise’ for individual effect observations that also takes causal principles into consideration  [F. Ribeiro et al., 2022].
@@ 4.4 Non-parametric Effect Estimation
The way both treatment assigments $A$, Eq.(18), and effects $Y$ are calculated, Eq.(18), in the present approach is related to other non-parametric causal effect estimators. Target learning is a semi-parametric approach that starts by specifying the mechanism by which individuals were assigned to different treatment levels. It does not assume random assignments. Specifically, taking the effect of a factor $a$ as the target parameter, it’s two main components are a treatment assignment mechanism $Q_{0}$, and an empirical distribution $P_{w}$. The first defines the probability distribution of receiving a particular treatment (e.g., the conditional probability of treatment given observed factors) and the second is the observed distribution of $W=X-\{a\}$ considering only those individuals who received the specific treatment (i.e., the distribution of the observed data for individuals who were treated according to a particular assignment mechanism). The targeted learning estimator $\hat{\Delta y(a)}_{n}$ is obtained by solving the following optimization problem:
$\hat{\Delta y(a)}_{n}=\arg\min_{\theta}\left\{\frac{1}{m-1}\sum_{w\in X-\{a\}}L\left(\Delta y(a),U_{0},P_{w}(a)\right)\right\}$
@@
Here, $L$ is the loss function over model-specific hyperparameters $\theta$ that quantifies the discrepancy between the true parameter $\Delta y(a)$, the data-generating distribution $Q_{0}$, and the empirical distribution $P_{w}$ induced by a specific treatment assignment. The optimization is typically performed using ensemble or black-box machine learning techniques. The optimization thus aims to find hyperparameter $\theta$ that minimizes the loss across all treatment assignment mechanisms, resulting in an estimator that is robust to different ways treatments may have been assigned in the non-experimental sample. The approach here has the same goal, but relies on a theoretical connection between treatment-background assignment mechanisms, in one hand, and their relation to generalizability and confoudness of effects on the other. The theory provides an answer to when specific effect observations can be said to have the later properties, conditional on a fixed mechanism (as identified by sample gap statistics). Our main goal is to demonstrate that under some mechanisms input samples are capable of generating effects that are externally valid. The same is true of parametric methods that employ black-box techniques across time to generate treatment-control pairs for effect estimation [Calonico et al., 2019, Abadie, 2021].
@@
In respect to outcome functions, Eq.(18), a general non-parametric strategy is to condition outcomes on all in-sample factors and take the difference in means for each stratum. This results in a treatment effect within each stratum of sample factors. One might now estimate the causal effect of treatment as the average over all strata of these strata-specific treatment effects. This approach is significantly more efficient than the unadjusted treatment effect estimators. That’s because the Law of Total Probabilities allows such methods to control (in-sample) confounding without the assumption that the effects are constant across strata (levels of confounders included in the model). If $U=\emptyset$, the g-formula  [Keil et al., 2014] implements this idea. For example, an outcome could be expressed as
$P(y=1)=\sum P(y=1|W=w)\times P(w)$ (21)
@@
which indicates that we are summing over each possible value of $W=X-\{a\}$, and $P(W=w)$ is the probability that $W$ takes on the value $w$ in the reference population and only $a$ is left to vary. This adjusted estimator of the effect of $a$ is a nonparametric maximum likelihood estimator for $Y$ for a single treatment and covariate. Drawbacks are the requirement that $U=\emptyset$ and the loss in performance in strata with small number of sample units. In Eq.(3), effects are conditioned on all their backgrounds (where order is also taken to matter). The strategy in Eq.(21) is often combined with conditional independence tests to identify plausible in-sample confounders. While we presently favor simplicity to highlight theoretical considerations and unobserved factors, elements of these approaches (such as further in-sample conditional independence tests of the g-formula and the supervised plug-in flexibility of targeted learning) could be further combined with the proposed.
@@ 6 Experiments
We consider simulations with increasing statistical complexity and in real-world applications. In both cases, we study three distinct prediction problems in held-out samples with $n$ observations. The tasks are to predict
@@
These correspond to out-of-sample prediction of outcomes, prediction of intervention effects, and marginal factor effect estimation (evaluated against the ground-truth). The first is the central problem in supervised Machine Learning, the second is related to effect estimation with counterfactual and matching estimators [Morgan and Winship, 2007], and the third to both causal effect estimation [Pearl, 2000], and supervised explainability  [Burkart and Huber, 2021]. Effect estimation is, in the present perspective, an extreme omitted-variable problem. The problem gets easier with the decrease of unobserved confounders and the proposed framework can also be used to study the performance of generic samples and models.
@@
The last number indicates that, to form a partial permutation, we select $d$ unique factors to fix and be organized in $\binom{m}{d}$ ways, each with $D_{m-d}$ possible disjoint orderings of the non-selected factors. Fig.3(a) shows the number of observed permutations present on increasingly larger samples (balanced). It illustrates how sample sizes affect the number of permutations asymptotically enumerable in them. Each (green) curve corresponds to the number of permutations in samples of size $n=\{200,350,500,1000\}$ (blue curves show the number of squares). The lower panel shows combinatorial limits (red) for the number of partial permutations and squares according to Eq.(23), $m=10$. The figure suggests, more specifically, necessary sample sizes $n$ such that the number of permutations $\Pi_{n}(Z)$ coincide with the limit of all partial permutations555since $q\leq 20$ in these illustrations all permutations in a sample can be enumerated with the approach in Appendix.I, for each number $d$ of fixation-points, $d\leq m$. The relationship between the DGP and enumerable permutations or their speed of enumeration is, for example, formulated in Eq.(2). We expect these non-parametric counts to impact biases and predictive performance, Eq.(3), irrespective of the DGP. The same is shown as histograms of square member counts, Fig.3(b). At $n{=}200$, no differences of size larger than $4$ are observed. A full square of size $10$ is completely observed only at $n{=}1024$ (bottom-right) in the balanced case, and $n\approx 20K$ in the unbalanced case. This was anticipated by Eq.(22) and follow from simple sample ordinal statistics. The first sample size is $\tilde{n}=2^{9+1}=1024$. The latter sample has rare factors, ${\rm I\kern-1.99997ptE}[min_{a\in X}(p(a))]=0.025$, which leads to an approximate sample requirement of $\tilde{n}=2^{9}/0.025=20480$. Similar sample size recommendations will be indicated for each of the increasingly complex simulated and real-world cases considered below.
@@ 6.1 Outcome Prediction
An understanding of the relationship among samples and estimators’ ACC and EV is necessary to devise accurate and maximally general models. To that end, we consider ACC (percentage of correctly classified cases in sample validation sections) as we increase $n$. For validation, we divide samples in two random sections: internal and external. The internal section is divided in typical training and validation subsections. We report accuracy of cross validation in the internal section (green) and external (red), with $4$ folds across cases. When ordered arbitrarily, or randomly, the internal-external division is inconsequential (as long as individual sections contain enough samples). We will define, however, alternative sample transversal orders, and consider how they impact learning performance. Starting with $n{=}1$, we train several algorithms with increasing $n$, and under different orders. With each unitary $n$ increase, an observation in the external section is transferred to the internal, and the problem of model and effect generalization becomes easier. Accordingly, a sample that can generalize its model and effects to external populations the earliest, for the same number of observations, can be said to have higher EV. We use three pre-specified XGBoost GBM (Gradient Boosting Machine) models, a grid of GLMs (Generalized Linear Model), a Random Forest (DRF), LASSO and Ridge regressions, five pre-specified GBMs, a near-default Deep Neural Net, an Extremely Randomized Forest (XRT), a random grid of XGBoost GBMs, a random grid of GBMs, and a random grid of Deep Neural Nets, as well as Stacked Ensembles with all previous models. Searched parameters are listed in Appendix.F.
@@
We start with the case of increasing EV (ACC in the external section). We expect a relationship between the number of squares in samples and their EV. According to Proposition.1 and Eq.(9), the square transversal order with fastest increase in ACC corresponds to a leftmost column first, $X_{(1)}$, transversal. This transversal is illustrated in Fig.4(d). We will consider other transversal orders, and their effects in respect to EV and CF, below. Performance of algorithms in samples with increasing sizes, under random and square transversals, are shown in Fig.4(a,b) for, respectively, the unbalanced and balanced cases. ACC Curves for 1000 simulation runs are shown for all algorithms (best parameter set), but bold curves mark the leader666we enumerate squares over the set $X$ of factors without the outcome to transverse samples, but repeat it for $X\cup\{y\}$ in each training fold; this leads to increase in ACC, without using data (outcomes) unavailable in typical validation protocols. (best algorithm and parameter-set combination). Error bars are shown only for the leader. For the random transversal case, ACC (both external and internal) remain at constant levels. Since sampling is random, algorithms minimize in this case the empirical, expected error in sample populations - which can reflect both population frequencies and sample selection biases. For the square case, there is a steep (linear) increase in EV. This indicates that maximizing the number of squares in data can lead to increases in EV. Learning in this case no longer minimizes the expected risk, but the risk for the combinatorial set of all subpopulations. In these cases, there is little difference in performance among different algorithms, with low variance across runs (bars), suggesting analytic limits and explanations. Recommended sample sizes for the many-squares asymptotic case, Eq.(22), are marked with a dotted line across figures (mean over runs).
@@
The figure also shows an alternative ’Travel-Salesman’ order (rightmost). It is obtained by solving a TSP [^7] Metric TSP  [Cormen, 2001](pg. 1029), in the worst case it generates solutions that are twice as long as the optimal tour (TSP calculation repeated in every simulation run). : sample units are cities, and their factor difference counts are distances. This order shares an important characteristic with square transversals: differences between units are kept small in this sampling order. The orders differ, however, in an important way: differences in the TSP are arbitrary and non-cyclic (following empirical sample frequencies). In this case, EV is steeply reducing, in contrast to square enumerations, Fig.4(a,b). This is not due to algorithms becoming sensitive to sample noise - as typical in overfitting. It portrays the expected behavior from algorithms when supplied with an increasingly complex population, generating models from specialized to general. The algorithmic stack includes regularized solutions (LASSO and Ridge regressions) and Target Learning (Sect. 4.4 Non-parametric Effect Estimation, Super-learner training, Appendix.F).

### v5

@@ 5.1 Randomization and Parametric Inference
We will use the notation often employed in Structurual Causal Models (SCM) to relate the previous concepts to research in parametric effect inference. In SCM models, variables are typically specified by an unknown function of a set of endogeneous variables and an exogeneous error. We have, similarly, specified DGPs $Z=\{X,U\}$ with a collection of exogenous factors $U$, in addition to the observed $X$. Factors in $U$ are unobserved and not affected by $X$, but can affect $X$. More specifically, a SCM specifies a collection of functions $f_{a}$ indexed by endogenous variables, $f=(f_{a}:a,b,c,...)$. The SCM also calls for the specification of the relation between the random variables $U$ and $X$. This is done through a set of functions $f=(f_{W},f_{a},f_{y})$, where $A$ is a random indicator for a treatment $a$, $W$ are the other (non-treated) in-sample factors, $W=X-\{a\}$, and $y$ is the outcome of interest. The values of $W$, $A$, and $Y$ are then deterministically assigned by $U$, according to  [van der Laan and Rose, 2011]
@@
Randomized controlled trials (RCTs) are one type of controlled experiment where subjects are randomized to receive a specific treatment. For example, when each subject is assigned to a treatment based on the flip of a coin. In this case, the differences between the outcomes of treatment and control groups become are due to treatment, as all other factors would be balanced (free of selection bias), up to random error. In an RCT, the treatment assignment mechanism, $P(A=1|W)$, is known (e.g., constant and $0.5$ as in the coin flip). This is different from non-experimental data where treatment, or treatment exposure, is not assigned. A RCT can be, nonetheless, written as in Eq.(16-17,19) in SCM notation,
@@
The only difference of Eq.(15) from Eq.(16-17,19) is the structural equation for $f_{a}$ in Eq.(17). The baseline variables $W$ play no role in the generation of $A$. These assumptions can also be illustrated with causal graphs, Fig.2(c). Since $a$ is randomized, this implies that $U_{a}$ is independent of $U_{y}$ and $U_{W}$, and, in this case, no arrows connect $U_{a}$ to $U_{y}$ and $U_{a}$ to $U_{w}$ in the graphical model specification. Eq.(LABEL:eq-rctthis) reviews the conditions on sample gap distributions proposed in this article and in Eq.(8).
With random assignment of backgrounds to fixed treatments, as the sample size $n$ approaches infinity, by the law of large numbers, the average characteristics of treatment and control groups, $P(Z-\{a\}<a)$ and $P(Z-\{a\}>a)$, converge to population averages. This can be written as
@@
With convergence, the potential bias introduced by correlations in observed-unobserved ($U-X$) is mitigated. On average, any unobserved or uncontrolled factors that could potentially bias effect estimates will be equally distributed between the two effect observation orders, as required by Principle (1). In a square, this is achieved by the stipulated requirements for the randomization of $Z-\{a\}$ factors, Eq.(14).
@@ 5.2 Non-parametric Effect Estimation
The way both treatment assigments $A$, Eq.(17), and effects $Y$, Eq.(19), are derived in the present approach are related to other non-parametric causal effect estimators. Targeted learning is a semi-parametric approach that starts by specifying the mechanism by which individuals were assigned to different treatments. It does not assume random assignments. Specifically, taking the effect of a factor $a$ as the target parameter, it’s two main components are a treatment assignment mechanism $Q$, and an empirical distribution $P_{w}$. The first defines the probability distribution of receiving a particular treatment (e.g., the conditional probability of treatment given observed factors) and the second is the observed distribution of $W=X-\{a\}$ considering only those individuals who received the specific treatment (i.e., the distribution of the observed data for individuals who were treated according to a particular mechanism). The targeted learning estimator $\hat{\Delta y(a)}_{n}$ is then obtained by solving the following optimization problem:
$\hat{\Delta y(a)}_{n}=\arg\min_{\theta}\left\{\frac{1}{m-1}\sum_{w\in X-\{a\}}L\left(\Delta y(a),Q,P_{w}(a)\right)\right\}$
@@
Here, $L$ is the loss function over hyperparameters $\theta$ that quantifies the discrepancy between the true parameter $\Delta y(a)$, the data-generating distribution $Q$, and the empirical distribution $P_{w}$ induced by a specific treatment assignment. The optimization is typically performed using ensemble or black-box machine learning techniques. The optimization thus aims to find hyperparameters $\theta$ that minimizes the loss across all treatment assignment mechanisms, resulting in an estimator that is robust to different ways treatments may have been assigned in the non-experimental sample. The approach here has the same goal, but relies on a theoretical connection between treatment-background assignments, in one hand, and their relation to generalizability and confoudness of effects on the other. The theory provides an answer to when specific effect observations can be said to have the later properties, conditional on a fixed mechanism (as identified by sample gap statistics and tests). We seek to demonstrate that under specific mechanisms samples contain effects that are externally valid. The approach holds a similar relationship to parametric methods that employ black-box techniques across time to generate treatment-control pairs for effect estimation [Calonico et al., 2019, Abadie, 2021].
@@
In respect to outcome functions, Eq.(19), a general non-parametric strategy is to condition outcomes on all in-sample factors and take their difference in means in each stratum. This leads to an estimate of the effect of treatment as the average over all strata. This approach is significantly more efficient than an unadjusted treatment effect estimate. That’s because the Law of Total Probabilities allows such methods to control (in-sample) confounding without the assumption that effects are constant across strata. If $q=0$, the g-formula  [Keil et al., 2014] implements this idea. For example, an outcome could be expressed as
$P(y=1)=\sum_{W=X-\{a\}}P(y=1|W=w)\times P(w)$ (21)
@@
which indicates that we are summing over each possible value of $W=X-\{a\}$, and $P(W=w)$ is the probability that $W$ takes on the value $w$ in the reference population when only $a$ is left to vary. This adjusted estimator of the effect of $a$ is a nonparametric maximum likelihood estimator for $Y$ for a single treatment  [Keil et al., 2014, van der Laan and Rose, 2011]. Drawbacks are the requirement that $q=0$ and the loss in performance in strata with small number of sample units. Analogously, in Eq.(3) effects are conditioned on all their backgrounds (where order is also taken to matter). The strategy in Eq.(21) is often combined with conditional independence tests to identify plausible in-sample confounders. While we presently favor simplicity to highlight theoretical considerations and unobserved confounding, elements of these approaches (such as further in-sample conditional independence tests of the g-formula and the supervised plug-in flexibility of targeted learning) could be further combined with the proposed.
@@ 7 Experiments
We consider performance of supervised, black-box explaining, and effect inference algorithms across a range of simulations and a real-world application. We thus study three distinct prediction problems in held-out samples with $n$ observations. The tasks are to predict
@@
These correspond to out-of-sample prediction of outcomes, prediction of intervention effects, and marginal factor effect estimation (evaluated against the ground-truth in simulations). The first is the central problem in supervised Machine Learning, the second is related to effect estimation with counterfactual and matching estimators [Morgan and Winship, 2007], and the third to both causal effect estimation [Pearl, 2000], and supervised explainability  [Burkart and Huber, 2021]. Effect estimation is, in the present perspective, an extreme omitted-variable problem. The problem becomes easier with decreases in the number $q$ of unobserved factors and the proposed framework can also be used to study the performance of generic samples and models.
@@
The last number indicates that, to form a partial permutation, we select $d$ unique factors to fix and be organized in $\binom{m}{d}$ ways, each with $D_{m-d}$ possible disjoint orderings of the non-selected factors. Fig.3(a) shows the number of observed permutations present on increasingly larger samples (balanced). Each (green) curve corresponds to the number of permutations in samples of size $n=\{200,350,500,1000\}$ (blue curves show the number of squares). The lower panel shows combinatorial limits (red) for the number of partial permutations and squares according to Eq.(23), $m=10$. The figure illustrates, in particular, the relationship between sample sizes $n$ and the subset of permutations $\Pi_{n}(Z)$ that effects for factors $Z$ can be asymptotically observed under222since $q\leq 20$ in these illustrations all permutations in a sample can be enumerated with the approach in Appendix.I, for each number $d$ of fixation-points, $d\leq m$. The relationship between the DGP and enumerable permutations or their speed of enumeration was, for example, formulated in Eq.(2) and illustrated in Fig.1(b). We expect these quantities to impact sample biases and predictive performance, Eq.(3), irrespective of the DGP. The same is shown as histograms of square member counts, Fig.3(b). At $n{=}200$, no differences of size larger than $4$ are observed. A full square of size $10$ is completely observed only at $n{=}1024$ (bottom-right) in the balanced case, and $n\approx 20K$ in the unbalanced case. This was anticipated by Eq.(22) and follow from simple sample ordinal statistics. The first sample size is $\tilde{n}=2^{9+1}=1024$. The latter sample has rare factors, ${\rm I\kern-1.99997ptE}[min_{a\in X}(p(a))]=0.025$, which leads to an approximate sample requirement of $\tilde{n}=2^{9}/0.025=20480$. Similar sample size recommendations will be indicated for each of the increasingly complex simulated and real-world cases considered below.
@@ 7.1 Outcome Prediction
An understanding of the relationship among samples and estimators’ ACC and EV is necessary to devise accurate and maximally general subsamples. To that end, we consider ACC (percentage of correctly classified cases in sample validation sections) as we increase $n$. To that end, we divide samples in two random sections: internal and external. The internal section is divided in typical training and validation subsections. We report accuracy of cross validation in the internal section (green) and external (red), with $4$ folds across cases. When ordered arbitrarily, or randomly, the internal-external division is inconsequential (as long as individual sections contain enough samples to cross algorithm-specific overfitting thresholds). We will define, however, alternative sample transversal orders, and consider how they impact learning performance. Starting with $n{=}1$, we train several algorithms with increasing $n$, and under different orders. With each unitary increase of $n$, an observation in the external section is transferred to the internal, and the problem of model and effect generalization becomes easier. Accordingly, a sample that can generalize its model and effects to external populations the earliest can be said to have higher EV. We use: three pre-specified XGBoost GBM (Gradient Boosting Machine) models, a grid of GLMs (Generalized Linear Model), a Random Forest (DRF), LASSO and Ridge regressions, five pre-specified GBMs, a near-default Deep Neural Net, an Extremely Randomized Forest (XRT), a random grid of XGBoost GBMs, a random grid of GBMs, and a random grid of Deep Neural Nets, as well as Stacked Ensembles with all previous models. Searched and optimized parameters in each case are listed in Appendix.F.
@@
We start with the case of increasing EV (i.e., ACC in the external section). We expect a relationship between the number of squares in samples and their EV. According to Proposition (1) and Eq.(9), the square transversal order with fastest increase in ACC corresponds to a leftmost-column first, $X_{(1)}$, transversal. This transversal is illustrated in Fig.4(d). We will consider other transversals, and their effects in respect to EV and CF, below. Performance of algorithms in samples with increasing sizes, under random and square transversals, are shown in Fig.4(a,b) for, respectively, the unbalanced and balanced cases. ACC Curves for 1000 simulation runs are shown for all algorithms (best parameter set), but bold curves mark the leader (best algorithm and parameter-set combination). Error bars are shown only for the leader. For the random transversal case, ACC (both external and internal) remain at constant levels. Since sampling is random, algorithms minimize in this case the empirical, expected error in sample populations - which can reflect both population frequencies and sample selection biases. For the square case, there is a steep (linear) increase in EV. This indicates that maximizing the number of squares in data can lead to increases in EV. In these cases, there is little difference in performance among different algorithms, with low variance across runs (bars), suggesting analytic limits and explanations. Recommended sample sizes for the many-squares asymptotic case, Eq.(22), are marked with a dotted line across figures (mean over runs).
@@
The figure also shows an alternative ’Travel-Salesman’ order (rightmost). It is obtained by solving a TSP [^3] Metric TSP  [Cormen, 2001](pg. 1029), in the worst case it generates solutions that are twice as long as the optimal tour (TSP calculation repeated in every simulation run). : sample units are cities, and their factor difference counts are distances. This order shares an important characteristic with square transversals: differences between units are kept small in this sampling order. The orders differ, however, in an important way: differences in the TSP are arbitrary and non-cyclic (following empirical sample frequencies). In this case, EV is steeply reducing, in contrast to square enumerations, Fig.4(a,b). This is not due to algorithms becoming sensitive to sample noise - as typical in overfitting. It portrays the expected behavior from algorithms when supplied with increasingly varying inputs, generating models from specialized to general. The algorithmic stack includes regularized solutions (LASSO and Ridge regressions) and Targeted Learning (Sect. 5.2 Non-parametric Effect Estimation, Super-learner training, Appendix.F).